# Flow Matching at Scale: A Machine Learning Framework for Efficient Large-Size Sampling of Many-Body Systems

### v3

@@ VIII.3 Learning Local Physics Instead of Global Statistics
    Directed Dynamics : The process is governed by an ODE defined by a learned vector field, This vector field acts as a time-dependent force that deterministically ”flows” any initial random state $\mathbf{x}_{0}$ at time $t=0$ to a structured, physical state $\mathbf{x}_{1}$ at time $t=1$. Unlike MCMC, this process is explicitly designed to change the probability distribution over time, moving it along a predefined path from noise to data.
@@
    Learning Local Rules of Transformation : The model is trained to minimize the discrepancy between its vector field and a target vector field derived from the MCMC data. Crucially, the model is not learning the global statistical properties of the small training systems. Instead, it is learning a function that dictates, for any given configuration $\mathbf{x}_{t}$ at any intermediate time $t$, the optimal infinitesimal change required to make it more ”physics-like.” For the XY model, this involves learning local rules, such as encouraging adjacent spins to align to lower the energy or resolving topological defects like vortices.
@@
    Distillation of Scale-Invariant Principles: The Hamiltonian of the XY model is defined by local, nearest-neighbor interactions. The physical principles governing low-energy configurations, the preference for smooth spin variations and the structure of vortices, are therefore local and independent of the overall system size. Flow Matching, by learning the vector field from local configuration data, distills these scale-invariant physical rules. Powered by the inductive biases embedded in CNN architectures, the learned vector field at specific site effectively only ”sees” the configuration of its local neighborhood, not the global lattice boundaries.
The origin of system-size generalization is that Flow Matching learns local, scale-invariant physics rather than global, size-specific statistics. In contrast to MCMC—which, for a fixed lattice (e.g., $32\times 32$), explores the entire equilibrium manifold without embedding knowledge that transfers directly to a larger lattice (e.g., $128\times 128$)—Flow Matching discovers a continuous representation of the local laws governing spin interactions. Because these laws are local and translationally invariant, the learned vector field effectively acts on small spin patches and is intrinsically agnostic to the global size of the system.
Consequently, applying the trained model to a larger lattice does not extrapolate global statistical properties; it simply deploys the same learned local rules over a broader domain. The model generates valid large-scale configurations because the underlying physics it has internalized is the same at all scales. In essence, the generative model did not learn what a small XY system looks like; it learned what the XY model is at a fundamental, local level, thereby enabling the construction of physically plausible configurations of arbitrary size. This success underscores a powerful principle in physics-informed machine learning: the most effective models are often those whose architectural inductive biases are deliberately aligned with the intrinsic locality and symmetries of the physical system being studied, a principle that has proven effective in a growing body of literature at the intersection of machine learning and physics [18, 33, 34, 38, 39, 24, 40].
@@ IX.1 Comparison of Normalized Flow and Flow Matching
Normalizing Flows construct this transformation as a sequence of invertible and differentiable mappings. This framework’s core requirement is that each transformation’s Jacobian determinant must be computationally tractable. This is essential because NFs rely on the change of variables formula to compute the target distribution’s exact likelihood, which is typically used to minimize the Kullback-Leibler (KL) divergence during training. This constraint imposes severe restrictions on the neural network architectures that can be employed. To ensure a tractable Jacobian, NFs often utilize specialized structures like coupling layers [41, 42], which partition the input dimensions into active and passive sets. Consequently, these models require carefully designed problem-specific masking mechanisms to ensure that all dimensions are updated over a sequence of layers. This architectural rigidity makes the implementation more complex and potentially unstable, and may also break the inherent symmetries of the modeled system [38, 39], a significant drawback in physical applications.
@@ IX.2 Generator Learning Framework
Learning the generator in Flow Matching methods, specifically the velocity field along the probability path connecting different distributions, provides numerous advantages. By explicitly modeling this generator through a neural network, the mapping between distributions becomes transparent and interpretable, circumventing the opaque and implicit representations typically encountered in conventional neural networks. This explicit representation significantly improves the flexibility and applicability of the neural network model. Suppose that we treat the neural network as the generator of the distribution mapping, it becomes feasible to employ diverse training strategies beyond Flow Matching alone, such as traditional maximum likelihood estimation or reverse KL divergence minimization in training Normalizing Flows (but need to solve the probability flow ODE during the training process), which can also be adapted to suit various problem contexts and specific research objectives. Overall, this approach aligns well with contemporary efforts in machine learning to build more interpretable, flexible, and versatile generative models, enabling clearer insight into the learned representations and improving the model’s adaptability across a broader range of applications.
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s reduced accuracy in predicting second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix C. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 10), which excels at learning the mean of the conditional velocity field but may under-represent its complete variance. This characteristic, in fact, constrains the direct application of this framework to physical problems where high-precision information about fluctuations, such as calculating critical exponents from specific heat or susceptibility, is paramount. Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, drastically reducing the simulation time required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
Another limitation is that the Flow Matching framework is designed for systems with continuous degrees of freedom, as in the XY model. Its extension to discrete models (e.g., Ising or Potts) is non-trivial, as the continuous probability flow is not directly applicable. There are two main paths to address this challenge. The first involves reformulating the discrete problem into an equivalent continuous one, for instance, using an auxiliary field method like the Hubbard-Stratonovich transformation, making the system tractable for our approach. A second, more direct path is to leverage generative models explicitly designed for discrete spaces. The development of frameworks such as Masked Diffusion [43] and more recent Discrete Flow Matching [44] marks a promising research frontier for applying these powerful sampling techniques to a broader class of physical systems. Indeed, several recent works have already started to explore such applications [45].
@@ X Conclusion and Outlook
In this work, we have introduced a generative model based on Flow Matching that offers a powerful alternative for the computational study of many-body systems. Our key finding is that a single network, by learning the scale-invariant, local physical rules of the Hamiltonian rather than the global statistics of a specific system size, can generate high-fidelity configurations across continuous temperatures and, crucially, for significantly larger lattices. This dual-generalization capability fundamentally alters the workflow for studying critical phenomena. It provides an efficient ”train-small, predict-large” strategy that can supplement or replace the traditional, brute-force approach of running separate, costly simulations for each system size. Our analysis shows that this approach provides a significant net computational advantage for large-scale studies, making robust finite-size scaling analysis more accessible.
The principles underpinning this method’s success—aligning the inductive biases of a CNN architecture with the locality inherent in physical Hamiltonians—suggest its applicability to a broader class of models. The framework is particularly well-suited for systems described by continuous fields on a lattice, opening promising avenues for its application in lattice field theories, such as scalar $\phi^{4}$ theory. Furthermore, its utility could extend to quantum many-body problems, especially those where quantum Monte Carlo methods map the system onto a classical statistical model with continuous variables. While our analysis highlights reduced accuracy for second-moment observables, this limitation opens a promising avenue for a powerful hybrid workflow. The core computational advantage is preserved by using the configurations generated by our Flow Matching model as high-quality, decorrelated initial states for a brief MCMC refinement. This synergistic approach, where FM acts as a potent accelerator and MCMC ensures rigorous convergence, paves the way for achieving unprecedented efficiency in high-precision tasks, such as accurately determining critical exponents. This establishes our methodology as an efficient sampling tool and a foundational component for next-generation, hybrid simulation strategies in statistical mechanics and beyond.
@@ Appendix A Details of Loss Function
In Eq (10) which we can show that the gradient is same as in Eq (9):
@@ Appendix B Training and Sampling Strategy
Experience has revealed that Flow Matching will have a relatively large error when the SNR is 1, so sampling time from a logit normal distribution in Eq. (30) with $\mu=0$ and $\sigma=1$ during the training phase would be a good choice.[46]
@@
The advantage of this time schedule is that one can flexibly adjust the mean and standard deviation to adapt to different sampling situations, and it is an effective means in practice.[46][47] In this work, the $(\mu,\sigma)$ are set to $(0,1)$ during training phase and (0.3,1)(-0.3,1) during inference phase.
All models in this work were trained using an identical hyperparameter configuration. We employed the AdamW optimizer [48] with the 1-cycle learning rate policy [49], setting the maximum learning rate to $2\times 10^{-4}$. Training was conducted with a batch size of 32, and we used 4 steps of gradient accumulation, resulting in an effective batch size of 128 for each parameter update.
@@ Appendix C More Physics Observables in Size Extrapolation
Figure 16: Size extrapolation of spin stiffness $\rho_{s}(T)$ for various system sizes ($L=16,64,128$), generated from a single FM model trained on $32\times 32$ data. The statistical noise visible across all curves, including the MCMC (MC) reference, is partially attributable to the modest ensemble size of 1500 configurations per temperature used for both training and comparison. While the FM model captures the qualitative behavior, a rigorous finite-size scaling analysis for the BKT transition requires locating the intersection of the stiffness curve with the universal line $2T/\pi$, according to the Nelson-Kosterlitz criterion [50]. As the figure shows, for extrapolated sizes ($L=64,128$), a noticeable deviation between the FM and MC results appears near the critical transition temperature where this intersection is expected. This discrepancy invalidates a scaling analysis based on the intersection point and justifies our choice to use the more robust susceptibility peak (Sec. V).
@@ Appendix D Neural Network Architecture
Our model employs a U-Net architecture, a design renowned for its efficacy in image-to-image tasks due to its hierarchical structure that captures multi-scale features [16]. The network is composed of an encoder path, a bottleneck, and a decoder path, with skip connections linking corresponding levels of the encoder and decoder to preserve high-resolution spatial details. To handle the conditional nature of our generation task, the model integrates time and temperature embeddings, which are processed and fused before modulating the network’s behavior at each block.
@@ D.1 Core Building Blocks
The fundamental component of our U-Net is the ”ResnetBlock” (RB). Each block consists of two main sequential operations: a spatial processing layer and a feedforward neural network. The input to each operation is first normalized using Root Mean Square Layer Normalization, a variant of Layer Normalization known for its computational efficiency and performance [51].
A key feature of the RB (see Fig. 17) is its adaptive conditioning mechanism, inspired by the Adaptive Layer Normalization (adaLN) used in Diffusion Transformers [37]. The conditional embedding, derived from time and temperature, is passed through a linear layer to generate shift and scale parameters. These parameters modulate the normalized feature maps within the block, allowing the network’s behavior to be dynamically controlled by the input conditions.
While the U-Net’s convolutional layers excel at learning local physical rules due to their inherent inductive biases of locality and translation equivariance, they are less effective at capturing the long-range correlations that emerge near critical points. To address this limitation, we replace the standard convolutional layers in RB with a specialized multi-head self-attention mechanism [52] in the network’s deeper, lower resolution levels. The structure is shown in Fig. 17, which is denoted as ”ResnetBlock (Attention)” (RBA) in the figure.
This attention mechanism incorporates a learnable 2D Rotary Position Embedding (RoPE), which has demonstrated a remarkable ability to handle multi-scale inputs and extrapolate across resolutions [53]. By applying rotational transformations to the query and key vectors based on their spatial coordinates, 2D RoPE effectively encodes relative positional information directly into the attention mechanism. This allows the model to capture complex, long-range dependencies between spin sites across the entire lattice, a crucial capability for modeling systems at or near a phase transition where correlation lengths diverge. The frequencies of the rotary embeddings are treated as learnable parameters, enabling the model to adapt the positional encoding to the specific physics of the system.

### v4

@@ VIII.3 Capturing Local Dynamics for Scalable Generation
    Directed Dynamics : The process is governed by an ODE defined by a learned vector field, This vector field acts as a time-dependent force that deterministically ”flows” any initial random state $\mathbf{x}_{0}$ at time $t=0$ to a structured, physical state $\mathbf{x}_{1}$ that is statistically consistent with the physical ensemble. at time $t=1$. Unlike MCMC, this process is explicitly designed to change the probability distribution over time, moving it along a predefined path from noise to data.
@@
    Approximating Local Dynamics of Transformation : The model is trained to minimize the discrepancy between its vector field and a target vector field derived from the MCMC data. Crucially, the model is not learning the global statistical properties of the small training systems. Instead, it is learning a function that dictates, for any given configuration $\mathbf{x}_{t}$ at any intermediate time $t$, the optimal infinitesimal change required to align the configuration with the target distribution. For the XY model, this involves learning local update tendencies, such as encouraging adjacent spins to align to lower the energy or resolving topological defects like vortices.
@@
    Distillation of Scale-Invariant Patterns: The Hamiltonian of the XY model is defined by local, nearest-neighbor interactions. The physical principles governing low-energy configurations, the preference for smooth spin variations and the structure of vortices, are therefore local and independent of the overall system size. Flow Matching, by learning the vector field from local configuration data, distills these scale-invariant interaction patterns. Powered by the inductive biases embedded in CNN architectures, the learned vector field at specific site effectively only ”sees” the configuration of its local neighborhood, not the global lattice boundaries.
The origin of system-size generalization lies in the distinction between learning global statistics and approximating local dynamics. Flow Matching, in our framework, captures the local conditional velocity field—effectively, the ”force” required to align spins locally based on their immediate environment. Because these local dynamics are translationally invariant and independent of the total system volume, they can be seamlessly applied to larger lattices.
Crucially, this explains how finite-size effects (FSE) are captured despite the model being size-agnostic. FSEs are not explicitly learned or memorized from the small training lattice. Instead, they emerge during the inference process as a consequence of applying the learned local dynamics on a manifold with larger spatial dimensions and periodic boundary conditions. The global correlation length is allowed to grow naturally until it is bounded by the new system size ($L$), thereby reproducing the correct scaling behavior ($T_{peak}(L)$) without requiring size-specific training data.
Consequently, applying the trained model to a larger lattice does not extrapolate global statistical properties; it simply deploys the same learned local rules over a broader domain. The model generates valid large-scale configurations because the underlying physics it has internalized is the same at all scales. In essence, it learned to approximate the local generative process of the XY model, thereby enabling the construction of physically plausible configurations of arbitrary size. This success underscores a powerful principle in physics-informed machine learning: the most effective models for locating phase transitions are often those whose architectural inductive biases are deliberately aligned with the intrinsic locality and symmetries of the physical system being studied, a principle that has proven effective in a growing body of literature at the intersection of machine learning and physics [Gerdes_2023LQF_equi_flow, Schuh:2025fky_Hubbard_Model_NF, Equivariant_flow_matching, PhysRevLett.125.121601, PhysRevD.103.074504, zhu2025physics_conditioned_DM, Hashimoto_2024].
@@ IX.1 Comparison of Normalized Flow and Flow Matching
Normalizing Flows construct this transformation as a sequence of invertible and differentiable mappings. This framework’s core requirement is that each transformation’s Jacobian determinant must be computationally tractable. This is essential because NFs rely on the change of variables formula to compute the target distribution’s exact likelihood, which is typically used to minimize the Kullback-Leibler (KL) divergence during training. This constraint imposes severe restrictions on the neural network architectures that can be employed. To ensure a tractable Jacobian, NFs often utilize specialized structures like coupling layers [real_nvp, nice], which partition the input dimensions into active and passive sets. Consequently, these models require carefully designed problem-specific masking mechanisms to ensure that all dimensions are updated over a sequence of layers. This architectural rigidity makes the implementation more complex and potentially unstable, and may also break the inherent symmetries of the modeled system [PhysRevLett.125.121601, PhysRevD.103.074504], a significant drawback in physical applications.
@@ IX.2 Generator Learning Framework
Learning the generator in Flow Matching methods, specifically the velocity field along the probability path connecting different distributions, provides numerous advantages. By explicitly modeling this generator through a neural network, the mapping between distributions becomes explicit and mathematically tractable, circumventing the implicit representations typically encountered in conventional neural networks. This explicit representation significantly improves the flexibility and applicability of the neural network model. Furthermore, framing the neural network as the generator of the distribution mapping, it becomes feasible to employ diverse training strategies beyond Flow Matching alone, such as traditional maximum likelihood estimation or reverse KL divergence minimization in training Normalizing Flows (but need to solve the probability flow ODE during the training process), which can also be adapted to suit various problem contexts and specific research objectives. Overall, this approach aligns well with contemporary efforts in machine learning to build more interpretable, flexible, and versatile generative models, enabling clearer insight into the learned representations and improving the model’s adaptability across a broader range of applications.
@@ IX.3 Limitations of Flow Matching
A significant limitation highlighted by our analysis is the model’s tendency to underestimate second-moment observables, which are directly related to physical fluctuations. This is evident in the increased MSE for susceptibility as system size grows (see Fig. 9) and, more strikingly, in the statistical noise observed in the extrapolated spin stiffness, as detailed in Appendix C. This behavior is likely rooted in the nature of the regression-based training objective (Eq. 11), which prioritizes the mean-field trend of the conditional velocity field, often at the expense of capturing the full variance. This characteristic, in fact, limits the utility of the generated ensemble for directly extracting critical exponents that rely on precise fluctuation magnitudes. However, as shown in Sec. V, the model remains highly effective for locating the critical temperature ($T_{c}$). Improving the model’s ability to capture the variance of the target distribution faithfully is therefore a critical challenge for future research. While potential avenues could include exploring alternative loss functions, a more immediate and practical solution is integrating our method into a powerful hybrid workflow. In this paradigm, the configurations generated by the FM model serve as high-quality, decorrelated initial states for a subsequent, brief refinement phase using traditional MCMC. A short MCMC chain can efficiently correct for any residual biases from the generative process, ensuring the final ensemble rigorously converges to the target Boltzmann distribution. Combining initial MCMC seeding, massive-scale FM sampling for interpolation and extrapolation, and a final MCMC refinement preserves the core computational advantage. The Flow Matching stage still acts as a potent accelerator, significantly reducing the computational cost required to obtain a large set of independent samples suitable for precision tasks like accurately determining critical exponents.
Another limitation is that the Flow Matching framework is designed for systems with continuous degrees of freedom, as in the XY model. Its extension to discrete models (e.g., Ising or Potts) is non-trivial, as the continuous probability flow is not directly applicable. There are two main paths to address this challenge. The first involves reformulating the discrete problem into an equivalent continuous one, for instance, using an auxiliary field method like the Hubbard-Stratonovich transformation, making the system tractable for our approach. A second, more direct path is to leverage generative models explicitly designed for discrete spaces. The development of frameworks such as Masked Diffusion [masked_diffusion] and more recent Discrete Flow Matching [gat2024discrete_flow_matching] marks a promising research frontier for applying these powerful sampling techniques to a broader class of physical systems. Indeed, several recent works have already started to explore such applications [tuo2025scalablemultitemperaturefreeenergy].
@@ X Conclusion and Outlook
In this work, we have introduced a generative model based on Flow Matching that offers an efficient framework for estimating critical properties in many-body systems. Our key finding is that a single network, by capturing the scale-invariant local interaction patterns of the Hamiltonian rather than the global statistics of a specific system size, can generate configurations that reflect essential critical correlations across continuous temperatures and, crucially, for significantly larger lattices. This dual-generalization capability fundamentally alters the workflow for studying critical phenomena. It provides an efficient ”train-small, predict-large” strategy that can complement the traditional, brute-force approach of running separate, costly simulations for each system size. Our analysis shows that this approach provides a significant net computational advantage for large-scale studies, making the identification of phase transitions via finite-size scaling analysis more accessible.
The principles underpinning this method’s success—aligning the inductive biases of a CNN architecture with the locality inherent in local interaction models suggest its applicability to a broader class of models. The framework is particularly well-suited for systems described by continuous fields on a lattice, opening promising avenues for its application in lattice field theories, such as scalar $\phi^{4}$ theory. Furthermore, its utility could extend to quantum many-body problems, especially those where quantum Monte Carlo methods map the system onto a classical statistical model with continuous variables. While our analysis highlights the tendency to underestimate second-moment observables, this limitation opens a promising avenue for a powerful hybrid workflow. The core computational advantage is preserved by using the configurations generated by our Flow Matching model as high-quality, decorrelated initial states for a brief MCMC refinement. This synergistic approach, where FM acts as a potent accelerator and MCMC ensures rigorous convergence, paves the way for achieving enhanced efficiency in high-precision tasks, such as accurately determining critical exponents. This establishes our methodology as an efficient sampling tool and a foundational component for next-generation, hybrid simulation strategies in statistical mechanics and beyond.
@@ Appendix A Details of Loss Function
We can show that the gradient of the conditional loss in Eq. (11) is identical to that of the original objective in Eq. (10):
@@ Appendix B Training and Sampling Strategy
Experience has revealed that Flow Matching will have a relatively large error when the SNR is 1, so sampling time from a logit normal distribution in Eq. (31) with $\mu=0$ and $\sigma=1$ during the training phase would be a good choice.[esser2024SD3]
@@
The advantage of this time schedule is that one can flexibly adjust the mean and standard deviation to adapt to different sampling situations, and it is an effective means in practice.[esser2024SD3][labs2025_flux1_kontext_flowmatching] In this work, the $(\mu,\sigma)$ are set to $(0,1)$ during training phase and (0.3,1)(-0.3,1) during inference phase.
All models in this work were trained using an identical hyperparameter configuration. We employed the AdamW optimizer [loshchilov2019decoupled_weight_decay] with the 1-cycle learning rate policy [smith2019superconvergence], setting the maximum learning rate to $2\times 10^{-4}$. Training was conducted with a batch size of 32, and we used 4 steps of gradient accumulation, resulting in an effective batch size of 128 for each parameter update.
@@ Appendix C More Physics Observables in Size Extrapolation
Figure 16: Size extrapolation of spin stiffness $\rho_{s}(T)$ for various system sizes ($L=16,64,128$), generated from a single FM model trained on $32\times 32$ data. The statistical noise visible across all curves, including the MCMC (MC) reference, is partially attributable to the modest ensemble size of 1500 configurations per temperature used for both training and comparison. While the FM model captures the qualitative behavior, a rigorous finite-size scaling analysis for the BKT transition requires locating the intersection of the stiffness curve with the universal line $2T/\pi$, according to the Nelson-Kosterlitz criterion [PhysRevLett.39.1201]. As the figure shows, for extrapolated sizes ($L=64,128$), a noticeable deviation between the FM and MC results appears near the critical transition temperature where this intersection is expected. This discrepancy invalidates a scaling analysis based on the intersection point and justifies our choice to use the more robust susceptibility peak (Sec. V).
@@ Appendix D Neural Network Architecture
Our model employs a U-Net architecture, a design renowned for its efficacy in image-to-image tasks due to its hierarchical structure that captures multi-scale features [ronneberger2015unet]. The network is composed of an encoder path, a bottleneck, and a decoder path, with skip connections linking corresponding levels of the encoder and decoder to preserve high-resolution spatial details. To handle the conditional nature of our generation task, the model integrates time and temperature embeddings, which are processed and fused before modulating the network’s behavior at each block.
@@ D.1 Core Building Blocks
The fundamental component of our U-Net is the ”ResnetBlock” (RB). Each block consists of two main sequential operations: a spatial processing layer and a feedforward neural network. The input to each operation is first normalized using Root Mean Square Layer Normalization, a variant of Layer Normalization known for its computational efficiency and performance [zhang2019RMSNorm].
A key feature of the RB (see Fig. 17) is its adaptive conditioning mechanism, inspired by the Adaptive Layer Normalization (adaLN) used in Diffusion Transformers [peebles2023DiT]. The conditional embedding, derived from time and temperature, is passed through a linear layer to generate shift and scale parameters. These parameters modulate the normalized feature maps within the block, allowing the network’s behavior to be dynamically controlled by the input conditions.
While the U-Net’s convolutional layers excel at learning local physical rules due to their inherent inductive biases of locality and translation equivariance, they are less effective at capturing the long-range correlations that emerge near critical points. To address this limitation, we replace the standard convolutional layers in RB with a specialized multi-head self-attention mechanism [vaswani2023attention_need] in the network’s deeper, lower resolution levels. The structure is shown in Fig. 17, which is denoted as ”ResnetBlock (Attention)” (RBA) in the figure.
This attention mechanism incorporates a learnable 2D Rotary Position Embedding (RoPE), which has demonstrated a remarkable ability to handle multi-scale inputs and extrapolate across resolutions [heo2024RoPE2d]. By applying rotational transformations to the query and key vectors based on their spatial coordinates, 2D RoPE effectively encodes relative positional information directly into the attention mechanism. This allows the model to capture complex, long-range dependencies between spin sites across the entire lattice, a crucial capability for modeling systems at or near a phase transition where correlation lengths diverge. The frequencies of the rotary embeddings are treated as learnable parameters, enabling the model to adapt the positional encoding to the specific physics of the system.